OpenStudy (anonymous):

Here Sacapuntas! (Ignore the problem) Choose the graph which matches the following system of equations. 5x + 2y = 2 3x - 3y = 18

OpenStudy (anonymous):

Ok. So let me look at the graphs again and I'll walk you through this.

OpenStudy (anonymous):

The first thing you want to do is get the system into a form you can graph. Standard form is good for solving, not so good for graphing. You remember how to solve for y, right?

OpenStudy (anonymous):

O.O No..

OpenStudy (anonymous):

You could possibly do the work first for all of them, then I could look over what you did.?

OpenStudy (anonymous):

I'll do the top one. Then you do the bottom one. :D

OpenStudy (anonymous):

BUT BUT BUT

OpenStudy (anonymous):

No, seriously, you have to get good at this. If you cannot solve for variables you will fail math forever. It's not that hard though, honest.

OpenStudy (anonymous):

Alright, I agree. How about this. Can I ask questions about your work?

OpenStudy (anonymous):

Of course!

OpenStudy (anonymous):

After I have said to myself... Hey, I know this! BUT.. I don't know that.

OpenStudy (anonymous):

Then I would ask that... ye

OpenStudy (anonymous):

\[5x + 2y = 2\]\[2y = 2 - 5x\]\[y = \frac{2 - 5x}{2}\]\[y = -\frac{5}{2}x + 1\]

OpenStudy (anonymous):

Ask if you don't understand!

OpenStudy (anonymous):

Alright, Now I am confused about how we may graph this. (I am looking at the graphs I sent you)

OpenStudy (anonymous):

You know \(y = mx + b\)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sorry, Yes! was getting food.

OpenStudy (anonymous):

Ok, so m is the slope and b is the y-intercept. From that we can make a graph!

OpenStudy (anonymous):

Teach me master >:)

OpenStudy (anonymous):

Ok, well, do you know what slope means?

OpenStudy (anonymous):

Means?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

Ok, well, mathematically, slope is just \(\frac{\Delta y}{\Delta x}\)

OpenStudy (anonymous):

Ok simple I guess

OpenStudy (anonymous):

It is. So, for example, a slope of \(\frac{5}{2}\) means for every 5 you go up, you go right 2.

OpenStudy (anonymous):

If the slope is negative, just go down instead of up. Or left instead of right. Whichever way is more convenient to think about. (You'll notice that reversing them both works the same way: for a slope of \(\frac{5}{2}\), every 5 you go down, you go left 2.)

OpenStudy (anonymous):

Ok so far so good

OpenStudy (anonymous):

Now, do you know what y-intercept means?

OpenStudy (anonymous):

I would guess that is referring to the Y intercept on a graph?

OpenStudy (anonymous):

Yes. But do you know what point that is? Like if my y-intercept is 2, what point is that?

OpenStudy (anonymous):

like up 2

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

So we know that point. And we also have a ratio that tells us "go to the right so many and then go up so many."

OpenStudy (anonymous):

In other words, that's all we need to graph a line. :D

OpenStudy (anonymous):

y=−5/2x+1?

OpenStudy (anonymous):

Yes. So start by going up 1. That's our y intercept. Now go to the left 2 and up 5. (You're going left because the slope is negative!) Or you can go down 5 and to the right 2. You'll get a different point but notice that all three points end up on the same line.

OpenStudy (anonymous):

Ok I am going to look at the graphs

OpenStudy (anonymous):

Wait was # Was this problem?

OpenStudy (anonymous):

It was the first one you gave me. So #1 I guess.

OpenStudy (anonymous):

Choose the graph which matches the following system of equations. 5x + 2y = 2 3x - 3y = 18

OpenStudy (anonymous):

Ok So we have one, now we need 3x - 3y = 18

OpenStudy (anonymous):

That's the one I'm going to let you do yourself! :D

OpenStudy (anonymous):

Idk :(

OpenStudy (anonymous):

I am sooooo bad

OpenStudy (anonymous):

Try! Follow the same method I used. You'll have a little time because I need to leave for a bit. I'll be back, though.

OpenStudy (anonymous):

B is what I got